OpenStudy (anonymous):

2 over 5 + 1 over 3 =

OpenStudy (anonymous):

11/15

OpenStudy (anonymous):

Do you mean (2/5)+(1/3)?

OpenStudy (karatechopper):

2/5+1/3 =11/15

OpenStudy (anonymous):

i did it but it was wrong help me lol

OpenStudy (anonymous):

You need to multiply the first fraction by 3/3 and the second fraction by 5/5 to get 6/15 + 5/15 = 11/15

OpenStudy (karatechopper):

ok so u have to make the fractions be commmon denominators

OpenStudy (anonymous):

ok